Project

General

Profile

Actions

Ops/Comm Scheduling meeting : 25/4/2017

Location : CR3 at 14:00 pm

Phone access : New phonebridge number
UK: 08082380274
US: 631.621.5253
Access code: 1779339517

Agenda

  • Actions
  • DC Update
  • LH2 Update
  • Tracker Issue
  • List of Hall activities in the shutdown (Colin)
  • Schedule for startup
    • Run plan discussion tomorrow
    • Proposed startup run : Thursday 4th May. May slip due to tracker cooldown.
    • Request readiness reports for run plan discussion and next week's meeting for straight track running from
      • Tracker
      • EMR/CKOV/KL
      • TOF
      • DAQ/Online
      • Controls and Monitoring
        • a email to Steve will be fine. Please include a readiness schedule

AOB:

Minutes

  • Actions
    • Ajit - No further development of Hanlet SMS system. Need to propose a basic messaging system still to alleviate need for 24Hr monitoring in person. - on hold
      • CWhyte to check with ABross/JBoehm that automated texting would be acceptable for removing requirement for continuous oversight while CC magnets powered.
      • We have the SMS to email gateway, including monitoring for e-mail responses, but a lot of work remains in driving this from the Alarm Handler.
    • Macwaters & Boehm - Arrange to install new Sumitomo compressor adsorbers - stands
    • Macwaters - Re-make dodgy compressor2 remote cable connector - stands
    • Macwaters - Prepare new F70 compressors for use. Remove old overheated CSW71 to R9. Speak to Sumitomo on April 5th regards how to repair. - in progress
    • Macwaters - Replace old vac sensors and new valve on cryo 1&2 - in progress
    • Ajit - Implement MOM/DC state machine override function for all systems; also need documentation and training for operators - documentation and training in progress
    • Macwaters - Report on status of FC psu repair (stands)
      • This needs to be cross-checked - e.g. Josef Boehm to check with Trevor. What is plan when it does return - will we want to test it in situ before the run?
    • Hall Manager - water system change-over procedure (stands)
    • EOverton - Status of miceraid5 disk cleanup (stands)
    • MUchida - investigate moving Tracker waveguides around to cure "hole". (superseded - see below)
    • MCourthold - show JBoehm how to use the Picologger for quench data. Craig can help. (on hold)
      • The Picologger has been disconnected. SBoyd to check with CWhyte whether we want it or not
    • JBoehm & Macwaters - get the West Wall compressors correctly labelled, as various have been re-purposed (stands)
    • CWhyte to contact DL to clarify their Hall Task List items
    • PFranchini to go through servers and turn off IPMI in BIOS done
  • New Actions
    • The Picologger has been disconnected. SBoyd to check with CWhyte whether we want it or not
    • HNebrensky to send BLOC-chasing email
  • DC Update:
    Activities concentrating on LH2 work:
  • LH2 Update
    The enlarged vent pipe has been installed. Mark Tucker working on vacuum. LH2 HAZOP in progress now. Leak-checking last week - 1 pressure relief valve taken out of use.
    It would seem daft not take physics data with Neon if the rest of MICE is ready, but need to find out what the plan for Neon test actually is: how full will the absorber be? (Limit of 1.8l in system) How long will the Neon test last (i.e. when are LH2 experts available to monitor)? It has been stated that the Neon test would only last a day.
    Also now want to do a quench test - can the LH2 absorber survive a quench of the FC - with Neon. Not clear where this fits in, as the FC will need time to recover.
  • Tracker Issue
    Cryostat 2: followed the official procedure for warm-up and cool-down. Have lost 12% of channels on one cassette and 27% on the other. Thought to be due to aging seals.
    Are getting hold of a sample bottle to be able to confirm presence of water: would then pump-purge to dry it out, and have 20% extra channels so there is some scope to re-arrange waveguides to re-cover some channels.
    So have a plan but won't know outcome until latter part of next week.
  • Schedule for startup
    • Experts have been sent readiness report request
    • Start-up delayed until 8th May itself.
      • Need to think about shifter training
    • Still have big gaps in BLOC schedule.
      • N.B. clash between expert-led run Monday-day and Henry in at midnight to put away target and beamstop.
    • Probably can't guarantee Diffuser operation in field until data-taking; as no window for test

Updated by Nebrensky, Henry over 6 years ago ยท 12 revisions